OpenStudy (anonymous):

If secθ=−5/4with π/2<θ<π , find the exact value of the other 5 trig functions of θ . If secθ=−5/4with π/2<θ<π , find the exact value of the other 5 trig functions of θ . @Mathematics

OpenStudy (mertsj):

You're in the second quadrant. The cosine is -4/5 since it is the reciprocal of the secant. So this is a 3-4-5 right triangle. The side adjacent is 4, the side opposite is 3 and the hypotenuse is 5 sin = 3/5 cos = -4/5 tan = -3/4 cot = -4/3 sec is given csc = 5/3

OpenStudy (anonymous):

does this" π/2<θ<π" really ,matter when solving the question?